# Towards Agentic OS: An LLM Agent Framework for Linux Schedulers

### v3

@@ Abstract
Operating system schedulers suffer from a fundamental semantic gap, where kernel policies fail to understand application-specific needs, leading to suboptimal performance. We introduce SchedCP, the first framework that enables fully autonomous Large Language Model (LLM) agents to safely and efficiently optimize Linux schedulers without human involvement. Our core insight is that the challenge is not merely to apply a better LLM, but to architect a decoupled control plane that separates the AI’s role of semantic reasoning ("what to optimize") from the system’s role of execution ("how to observe and act"). Implemented as Model Context Protocol(MCP) server, SchedCP provides a stable interface with three key services: a Workload Analysis Engine, an evolving Scheduler Policy Repository, and an Execution Verifier that validates all AI-generated code and configure before deployment with static and dynamic analysis. We demonstrate this architecture’s power with sched-agent, a multi-agent system that autonomously analyzes workloads, synthesizes custom eBPF scheduling policies, and deploys them via the sched_ext infrastructure. Our evaluation shows that SchedCP achieves up to 1.79x performance improvement and 13x cost reduction compared to naive agentic approaches, all while maintaining high success rate. The code will be open-sourced.
@@ 1 Introduction
Prior reinforcement learning-based schedulers mao2019decima ; qiu2020firm lack semantic understanding of workloads. While LLMs openai2023gpt4 ; anthropic2024claude and agent frameworks autogen ; geminicli ; claudecode ; qian2024chatdev ; hong2023metagpt excel at code generation, naively applying them to scheduler development proves impractical: our experiments show generating a basic scheduler takes 33 minutes, costs \$6, and often degrades performance. The gap remains: existing methods lack semantic understanding, while LLMs lack the scaffolding for safe, efficient, and reliable systems integration despite prior LLM-eBPF work kgent .
We introduce a decoupled architecture with two components. First, SchedCP is a control plane framework providing safe AI-kernel interfaces with profiling, tracing and validation tools, allowing LLMs to customize the kernel scheduler with eBPF. It exposes kernel scheduling features via Model Context Protocol (MCP) through three core services: Workload Analysis Engine, Scheduler Policy Repository, and Execution Verifier. Second, sched-agent is the first autonomous multi-agent system that decomposes scheduler optimization into four specialized agents (Observation, Planning, Execution, Learning), demonstrating how LLMs can bridge the semantic gap between application requirements and kernel scheduling policies. This separation allows SchedCP to provide a generalizable framework for any AI agent, while sched-agent demonstrates semantic workload analysis and policy generation. Our evaluation shows sched-agent achieves up to 1.79× performance on kernel compilation, 2.11× P99 latency improvement and 1.60× throughput gain on schbench, 20% latency reduction for batch workloads, and 13× cost reduction.
@@ 2 Motivation
Linux scheduler optimization faces three barriers. First, a domain knowledge gap exists between developers and users: DevOps engineers lack insight into workload characteristics (latency-sensitive vs. throughput-oriented), while edge/personal device users lack both kernel optimization expertise and understanding of application-specific targets. Second, scheduler development requires mastering kernel programming, limiting innovation to few experts. Third, modern workloads exhibit complex dynamics: web traffic varies by orders of magnitude daily, build system parallelism changes with dependencies. Prior RL-based schedulers mao2019decima ; qiu2020firm ; zhang2024mrsch ; mao2019park require extensive training per workload type, lack semantic understanding to transfer across workloads, and cannot generate code. LLMs uniquely bridge these gaps by: (1) understanding natural language and source code semantics without task-specific training, (2) synthesizing correct eBPF schedulers based on automatic workload characterization, (3) reasoning about performance trade-offs and system constraints, (4) LLM can operate in the control plane to generate optimized code that runs natively with negligible runtime overhead, unlike traditional ML models that would cause unacceptable inference latency in the scheduler hot path.
We tested Claude Codeclaudecode , the state of the art LLM agent, with "write a FIFO scheduler in eBPF" from an empty folder, with all permissions and bash access. Of three attempts, only one succeeded. The second attempt produced pseudo-code after 6 minutes trying, and the third generated a scheduler tracer instead after 8 minutes development. The successful generation required 33 minutes, 221 LLM API calls, and 15+ iterations, costing \$6 (vs. 5 minutes typically for an expert developer). The generated code, for some workloads, exhibited poor quality with excessive overhead, performing worse than EEVDF. The agent required root access, could crash the system during testing, and lacked fallback mechanisms, which also raises safety concerns. These experiments reveal three critical challenges: Performance, ensuring AI schedulers outperform existing ones; Safety, preventing crashes, lockups, and starvation while minimizing privileges; and Efficiency, reducing the 33-minute, \$6 generation cost for practical deployment.
@@ 3 The SchedCP Framework Design and Implementation
The Observation Agent builds Workload Profiles by querying the Workload Analysis Engine strategically, starting with high-level summaries from process name, commands then requesting deeper profiling (perf stat, top) based on findings, synthesizing data into natural language descriptions and optimization goals while managing cost-precision tradeoffs. For kernel compilation, it produces profiles like “CPU-intensive parallel compilation with short-lived processes, inter-process dependencies, targeting makespan minimization.” The Planning Agent transforms profiles into optimization strategies via the Scheduler Policy Repository, following a decision hierarchy: configuring existing schedulers, generating patches, or composing new schedulers from primitives. The Execution Agent manages development, validation and deployment by synthesizing code artifacts, submitting to the Execution Verifier, interpreting results to refine code or fix logic issues. The Learning Agent completes the ICRL loop by analyzing deployment outcomes (e.g., 45% makespan reduction), enabling in-session adaptation and updating the repository with refined metrics, deployment contexts, and documented antipatterns.
@@ 5 Preliminary Evaluation
We validate SchedCP’s effectiveness through four research questions: configuring existing schedulers (RQ1), generating new schedulers for specific workloads (RQ2), cost and efficiency of scheduler generation (RQ3), and iterative refinement improvements (RQ4). Evaluation uses two machines: 86-core Intel Xeon 6787P with 758GB RAM running Linux 6.14, and 8-core Intel Core Ultra 7 258V with 30GB RAM running Linux 6.13. Agents use Claude Code (Opus 4), testing each case three times and averaging results. All experiments successfully create working custom scheduler configurations or eBPF programs. Future evaluation requires a complete benchmark.

### v4

@@ Abstract
Operating system schedulers suffer from a fundamental semantic gap, where kernel policies fail to understand application-specific needs, leading to suboptimal performance. We introduce SchedCP, the first framework that enables fully autonomous Large Language Model (LLM) agents to safely and efficiently optimize Linux schedulers without human involvement. Our core insight is that the challenge is not merely to apply a better LLM, but to architect a decoupled control plane that separates the AI’s role of semantic reasoning ("what to optimize") from the system’s role of execution ("how to observe and act"), thereby separating the optimization problem into two stages: goal-inference and policy-synthesis. Implemented as Model Context Protocol(MCP) server, SchedCP provides a stable interface with three key services: a Workload Analysis Engine, an evolving Scheduler Policy Repository, and an Execution Verifier that validates all AI-generated code and configurations before deployment with static and dynamic analysis. We demonstrate this architecture’s power with sched-agent, a multi-agent system that autonomously analyzes workloads, synthesizes custom eBPF scheduling policies, and deploys them via the sched_ext infrastructure. Our evaluation shows that SchedCP achieves up to 1.79x performance improvement and 13x cost reduction compared to naive agentic approaches, all while maintaining high success rate. The code is in https://github.com/eunomia-bpf/schedcp.
@@ 1 Introduction
Prior reinforcement learning-based schedulers mao2019decima ; qiu2020firm lack semantic understanding of workloads, are often limited to tweaking configurations within a problem space predefined by human engineers, preventing fully automatic system optimization. While LLMs openai2023gpt4 ; anthropic2024claude and agent frameworks autogen ; geminicli ; claudecode ; qian2024chatdev ; hong2023metagpt excel at code generation, naively applying them to scheduler development proves impractical: our experiments show generating a basic scheduler takes 33 minutes, costs \$6, and often degrades performance. The gap remains: existing methods lack semantic understanding, while LLMs lack the scaffolding for safe, efficient, and reliable systems integration despite prior LLM-eBPF work kgent .
Our work enables a system to drive its own optimization by decomposing the problem into two stages: a goal-inference stage that uncovers optimization goals and constraints from workloads, and a policy-synthesis stage that compiles them into eBPF scheduler policies. This approach is embodied in a decoupled architecture with two components. First, SchedCP is a control plane framework providing safe AI-kernel interfaces with profiling, tracing and validation tools, allowing LLMs to customize the kernel scheduler with eBPF. It exposes kernel scheduling features via Model Context Protocol (MCP) through three core services: Workload Analysis Engine, Scheduler Policy Repository, and Execution Verifier. Second, sched-agent is the first autonomous multi-agent system that decomposes scheduler optimization into four specialized agents (Observation, Planning, Execution, Learning), demonstrating how LLMs can bridge the semantic gap between application requirements and kernel scheduling policies. This separation allows SchedCP to provide a generalizable framework for any AI agent, while sched-agent demonstrates semantic workload analysis and policy generation. Our evaluation shows sched-agent achieves up to 1.79× performance on kernel compilation, 2.11× P99 latency improvement and 1.60× throughput gain on schbench, 20% latency reduction for batch workloads, and 13× cost reduction.
@@ 2 Motivation
Linux scheduler optimization faces three barriers. First, a domain knowledge gap exists between developers and users: DevOps engineers lack insight into workload characteristics (latency-sensitive vs. throughput-oriented), while edge/personal device users lack both kernel optimization expertise and understanding of application-specific targets. Second, scheduler development requires mastering kernel programming, limiting innovation to a few experts. Third, modern workloads exhibit complex dynamics: web traffic varies by orders of magnitude daily, build system parallelism changes with dependencies. Prior RL-based schedulers mao2019decima ; qiu2020firm ; zhang2024mrsch ; mao2019park require extensive training per workload type, lack semantic understanding to transfer across workloads, and only tweak configurations after engineers have already defined the entire problem space: selecting features, specifying knobs, and writing objective functions. LLMs uniquely bridge these gaps by: (1) using tools to dynamically explore diverse workloads to uncover application intent and structure, which is difficult to capture with conventional static or dynamic analysis; (2) applying a broad, pre-trained understanding of source code semantics to reason about performance trade-offs and structural dependencies, such as data-dependency hints; (3) synthesizing correct eBPF schedulers based on this analysis; and (4) operating in the control plane to generate optimized code that runs natively with negligible runtime overhead, unlike traditional ML models that would cause unacceptable inference latency in the scheduler hot path.
We tested Claude Codeclaudecode , the state-of-the-art LLM agent, with "write a FIFO scheduler in eBPF" from an empty folder, with all permissions and bash access. Of three attempts, only one succeeded. The second attempt produced pseudo-code after 6 minutes trying, and the third generated a scheduler tracer instead after 8 minutes of development. The successful generation required 33 minutes, 221 LLM API calls, and 15+ iterations, costing \$6 (vs. 5 minutes typically for an expert developer). The generated code, for some workloads, exhibited poor quality with excessive overhead, performing worse than EEVDF. The agent required root access, could crash the system during testing, and lacked fallback mechanisms, which also raises safety concerns. These experiments reveal three critical challenges: Performance, ensuring AI schedulers outperform existing ones; Safety, preventing crashes, lockups, and starvation while minimizing privileges; and Efficiency, reducing the 33-minute, \$6 generation cost for practical deployment.
@@ 3 The SchedCP Framework Design and Implementation
The Observation Agent builds Workload Profiles by querying the Workload Analysis Engine strategically, starting with high-level summaries from process name and commands then requesting deeper profiling (perf stat, top) based on findings, synthesizing data into natural language descriptions and optimization goals while managing cost-precision tradeoffs. For kernel compilation, it produces profiles like “CPU-intensive parallel compilation with short-lived processes, inter-process dependencies, targeting makespan minimization.” The Planning Agent transforms profiles into optimization strategies via the Scheduler Policy Repository, following a decision hierarchy: configuring existing schedulers, generating patches, or composing new schedulers from primitives. The Execution Agent manages development, validation and deployment by synthesizing code artifacts, submitting to the Execution Verifier, interpreting results to refine code or fix logic issues. The Learning Agent completes the ICRL loop by analyzing deployment outcomes (e.g., 45% makespan reduction), enabling in-session adaptation and updating the repository with refined metrics, deployment contexts, and documented antipatterns.
@@ 5 Preliminary Evaluation
We validate SchedCP’s effectiveness through four research questions: configuring existing schedulers (RQ1), generating new schedulers for specific workloads (RQ2), cost and efficiency of scheduler generation (RQ3), and iterative refinement improvements (RQ4). Evaluation uses two machines: 86-core Intel Xeon 6787P with 758GB RAM running Linux 6.14, and 8-core Intel Core Ultra 7 258V with 30GB RAM running Linux 6.13. Agents use Claude Code (Opus 4), testing each case three times and averaging results. All experiments successfully created working custom scheduler configurations or eBPF programs. Future evaluation requires a complete benchmark.